# The Pound-Drever-Hall Method for Superconducting-Qubit Readout

### v1

@@ Abstract
The highest performance optical-frequency measurements often rely on self phase-referenced techniques, such as the Pound-Drever-Hall (PDH) method Drever et al. (1983); Black (2001) for locking lasers and cavities. At lower frequency, and longer wavelength, the path-length stability requirements for phase measurements are less stringent, and many measurement tasks, such as superconducting-qubit readout, are simply carried out by comparing a measurement signal to an external phase reference; however, phase drift remains a serious technical challenge. Here we show that the intrinsically phase-stable PDH technique that originated in the optical frequency domain, and has been used for fast characterization of microwave resonators Lindström et al. (2011); de Graaf et al. (2014), can be transported to the quantum microwave regime, where it gives rise to improvements for superconducting qubit readout, in terms of long-time stability, scaling complexity, and signal amplification.
@@
where is the qubit-dependent phase, and are carrier and LO phase contributions, and is the phase accrued along the propagation path. Generator-induced errors depend on the quality of the source and its reference lock, while path-length fluctuations on signals at (e.g., from thermal drifts) remain a dominant source of phase instability even with high-performance references. The fluctuations resulting from these limitations are illustrated in the IQ-plane representation of qubit readout shown in Fig. 1(a). While such fluctuations can, in principle, be fixed through advanced instrumentation and rolling calibration, these solutions are expensive, complex, and delicate to set up, thereby posing a scaling challenge for large-scale quantum systems.
To address these limitations, we explore an alternative approach for detecting qubit-state-dependent frequency shifts on the resonator based on the PDH method. The readout signal in this multi-tone interrogation scheme consists of three copropagating tones: the carrier (), the upper sideband (), and lower sideband (). For phase stability, these three tones are produced from a single source, usually via phase modulation at , and the PDH readout signal, derived from the beat between the carrier and sidebands, takes the form
@@ Acknowledgements.
This project was supported by ARL Grants Nos. W911NF-24-2-0107, W911NF-19-2-0181, and W911NF-17-S-0003. IA, WCL, and AK received additional support from NSF Grant No. PHY2047732, the Sloan Fellowship Program, AFOSR Grant No. Grant No. FA9550-21-1-0129, the NSF QLCI grant OMA-2120757, and the University of Maryland. We thank Jun Ye, Martin Ritter, and Nathan Schine for helpful comments and discussions, and we thank Jim Phillips and the McRae Group for helpful conversations regarding fast characterization and Q-factor measurements of superconducting resonators using PDH.References

### v2

@@ Abstract
The highest performance optical-frequency measurements often rely on self phase-referenced techniques, such as the Pound-Drever-Hall (PDH) method Drever et al. (1983); Black (2001) for locking lasers and cavities. At lower frequency, and longer wavelength, the path-length stability requirements for phase measurements are less stringent, and many measurement tasks, such as superconducting-qubit readout, are simply carried out by comparing a measurement signal to an external phase reference; however, phase drift remains a serious technical challenge. Here we show that the intrinsically phase-stable PDH technique that originated in the optical frequency domain, and has been used for fast characterization of microwave resonators Lindström et al. (2011); de Graaf et al. (2014); Rath et al. (2025), can be transported to the quantum microwave regime, where it gives rise to improvements for superconducting qubit readout, in terms of long-time stability, scaling complexity, and signal amplification.
@@
where is the qubit-dependent phase, and are carrier and LO phase contributions, and is the phase accrued along the propagation path. Fluctuations in depend on the quality of the sources and reference locks, and can be eliminated by modulating a single generator to produce both the measurement and LO tones. However, path-length fluctuations on signals at (e.g., from thermal drifts) remain a source of phase instability even with high-performance references. The resulting phase fluctuations are illustrated in the IQ-plane representation of qubit readout shown in Fig. 1(a). While such fluctuations can be removed through rolling calibration, a natively insensitive method should be more scalable.
Here we explore an approach for detecting qubit-state-dependent frequency shifts on the resonator based on the PDH method. The readout signal in this multi-tone interrogation scheme consists of three copropagating tones: the carrier (), the upper sideband (), and lower sideband (). For phase stability, these three tones are produced from a single source, usually via phase modulation at , and the PDH readout signal, derived from the beat between the carrier and sidebands, takes the form
@@ Acknowledgements.
This project was supported by ARL Grants Nos. W911NF-24-2-0107, W911NF-19-2-0181, and W911NF-17-S-0003. IA, WCL, and AK received additional support from NSF Grant No. PHY2047732, the Sloan Fellowship Program, AFOSR Grant No. Grant No. FA9550-21-1-0129, the NSF QLCI grant OMA-2120757, and the University of Maryland. Research at UMD was also supported by the Maryland Quantum Materials Center. We thank Jun Ye, Martin Ritter, and Nathan Schine for helpful comments and discussions, and we thank Jim Phillips and the McRae Group for helpful conversations regarding fast characterization and Q-factor measurements of superconducting resonators using PDH.References
@@
- Rath et al. (2025) P. K. Rath, J. D. Philips, T. Yoon, K. R. Shirer, A. Fereidouni, and J. Pollanen, arXiv preprint arXiv:2512.06289 (2025).